OpenStudy (anonymous):

Sheila wants to deposit $2,145 into savings accounts at three different banks: Bank of the US, 7/4 Bank, and Catch Bank. She will deposit five times as much into Catch Bank as 7/4 Bank because they offer a higher interest rate. She also expects the Bank of the US deposit to be 30% of the total of the other two deposits. How much will Sheila deposit into the Bank of the US savings account?

OpenStudy (campbell_st):

let Bank 7/4 = x deposited Catch Bank = 5x deposited Total deposits so far = 6x Bank US = 30% of 6x or \[0.3 \times 6x = 1.8x\] = 1.8x deposited Total deposited = 7.8x which equals the money available to Sheila. then 7.8x = 2145 divide by 7.8 to find the value of x x = $275 Bank 7/4 = $275 Catch Bank = 275 x 5 = $1375 Bank US = 1.8 x 275 = $495

OpenStudy (anonymous):

thank you so much